OpenStudy (anonymous):

A particle that moves along a straight line has velocity (see below) meters per second after t seconds. How many meters will it travel during the first t seconds?

OpenStudy (anonymous):

\[v(t)=t^2e ^{-3t}\]

OpenStudy (anonymous):

integrate

OpenStudy (anonymous):

I know that, but I might need a little help on the steps.

OpenStudy (rogue):

\[Displacement = \int\limits_{0}^{t} t ^{2} e ^{-3t}dt\]

OpenStudy (anonymous):

Remember that \[v(t)=\frac{dx(t)}{dt}\]

OpenStudy (anonymous):

If you integrate like kristal said: \[x(t)=\int_0^tv(t)dt\]

OpenStudy (rogue):

\[Distance = \int\limits_{0}^{t}\left| t ^{2} e ^{-3t} \right| dt\]

OpenStudy (anonymous):

Do you need help to solve that integral?

OpenStudy (anonymous):

I do, I know it's integration by parts. But I don't know which to make f(x) or g(x).

OpenStudy (rogue):

\[\mu = t ^{2}, d \upsilon = e ^{-3t}\]

OpenStudy (rogue):

I think, following LIPET.